OpenStudy (barrelracer011):

HELP WILL MEDAL!!! Which of the following shows the correct steps to find the value of 8 to the power of 1 over 3?

OpenStudy (barrelracer011):

8 to the power of 1 over 3 equals 4 to the power of 3 whole to the power of 1 over 3 equals 4 to the power of 3 multiplied by 1 over 3 equals 4 8 to the power of 1 over 3 equals 2 to the power of 3 whole to the power of 1 over 3 equals 2 to the power of 3 multiplied by 1 over 3 equals 2 8 to the power of 1 over 3 equals 2 to the power of 4 whole to the power of 1 over 3 equals 2 to the power of 4 multiplied by 1 over 3 equals 16 16 to the power of 1 over 4 equals 8 to the power of 2 whole to the power of 1 over 4 equals 8 to the power of 2 multiplied by 1 over 4 equals 8

OpenStudy (anonymous):

@aum @Marina17 @robtobey can you answer this?

OpenStudy (anonymous):

\[\sqrt[3]{8}=\sqrt[3]{2^3}=2^{\frac{3}{3}}=2 \]

OpenStudy (aum):

\(\Large (8)^{\frac 13}= (2^3)^{\frac 13} = 2^{3~*~ \frac 13} = 2^1 = 2\)